OpenStudy (anonymous):

(a) What is [- sqrt(144)]? (b) What is [sqrt(- 144)]? I think it is -12, but donot know (b)

OpenStudy (anonymous):

\[\sqrt{-144}=12i\] as it becomes imaginary if you take the root of a negative number.

OpenStudy (anonymous):

b) is 12i where i is iota ...but its a complex no.

OpenStudy (anonymous):

[- sqrt(144)] = -12

OpenStudy (anonymous):

[sqrt(- 144)] = 12 i

OpenStudy (anonymous):

the first one is real but the second is imaginary

OpenStudy (anonymous):

So [sqrt(- 144)] must not be a real number?

OpenStudy (anonymous):

\[-\sqrt{144}=-12\]\[\sqrt{-144}=12i\]

OpenStudy (anonymous):

yes every sqrt(-number) = imaginary number

OpenStudy (anonymous):

if you will get the nth root of a number wherein n is even,the answer is an imaginary number

OpenStudy (anonymous):

given that the number inside the radical is negative